OpenStudy (anonymous):

Given a and 1/2 are the roots of the quadratic equation 2x^2+px-3=0, find the values of a and p. see whether my working is correct. @ganeshie8

OpenStudy (anonymous):

\[2(\frac{ 1 }{ 2 })^2 + p(\frac{ 1 }{ 2 })-3=0\]\[\frac{ 1 }{ 2 }+\frac{ p }{ 2 }-3=0\]\[\frac{ p }{ 2 }=\frac{ 5 }{ 2 }\]\[2p=10\]\[p=5\]

OpenStudy (anonymous):

\[a=-\frac{ 3 }{ 2 }\div \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[a=-3\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent! looks perfect!

ganeshie8 (ganeshie8):

there is another way to solve this problem if you know the sum of roots and product of roots stuff

ganeshie8 (ganeshie8):

2x^2+px-3=0 a and 1/2 are roots product of roots = a*1/2 = -3/2 so a = -3 sum of roots = -3 + 1/2 = -p/2 so p = 5

ganeshie8 (ganeshie8):

Oh I see you're already using product of roots to solve a :)

OpenStudy (anonymous):

this working is faster and neat :) thnx @ganeshie8

ganeshie8 (ganeshie8):

yw:)